Suppose all individuals are identical, and their monthly demand for Internet access from a certain leading

provider can be represented as p=5-(1/2)q where p is price in $ per hour and q is hours per month. The

firm faces a constant marginal cost of $1. The profit-maximizing two-part tariff yields total revenue of

A)$24.

B)$40.

C)$32.

D)$16.
Answer #1

Answer

The correct answer is (A) $24

Profit Maximizing Rule Under Two Part Pricing :

In order to maximize profit under two part pricing a firm should charge Price(p) equal to its marginal cost(MC) and produce that quantity at which Price = Marginal Cost. Also it should charge Fixed fee(or entry fee) equal to consumer surplus.

p = MC => 5-(1/2)q = 1

=> q = 8

Consumer surplus = (1/2)*base*height

Vertical intercept = 5-(1/2)*0 = 5 hence vertical intercept = (0 , 5)

Entry fee = consumer surplus = (1/2)*8*(5 - 1) = 16

Hence Total Revenue = Entry fee + p*q

= 16 + 1*8

= $24

Hence, the correct answer is (A) $24